OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 9 inches and a height of 4 inches. Pyramid B has a volume of 8,100 in3. How many times bigger is the volume of pyramid B than pyramid A?

OpenStudy (anonymous):

15 32 54 75

OpenStudy (anonymous):

@phi Completely confused

jhonyy9 (jhonyy9):

the volum of a pyramid you can calcule using formula V = area of base *height /3 then volum A = 9*9*4 /3 = 3*9*4 = 108 in^3 then now just divide the volum of B by volum of A

OpenStudy (phi):

They want you to figure out the volume of pyramid A Pyramid A is a square pyramid with a base side length of 9 inches and a height of 4 inches. volume is area of base * height then divide by 3 the base is a square with side = 9. area of a square is side * side can you do that ?

OpenStudy (anonymous):

Is the answer 75?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

How many times bigger is the volume of pyramid B than pyramid A? you can test it: 75 * vol A = vol B ? is this true?

OpenStudy (anonymous):

Yes :)